OpenStudy (anonymous):

Can I get some help solving this Discrete-Time Dynamical System? The updating function is f(x)=.75x=2 and I've worked it out to be f(5)=.75^5(16)+2(.75^4)+2(.75^3)+2(.75^2)+2(.75)+2 But I can't figure out how to solve that to be the full solution. Any thoughts?

OpenStudy (anonymous):

\(m_t=0.75m_{t-1}+2\) Above is our difference equation, but to solve, we need more information... what is the initial value(s)?

OpenStudy (anonymous):

\[f(5)=.75^5(16)+2(.75^4)+2(.75^3)+2(.75^2)+2(.75)+2\] is what it really looks like. I just figured out the equation editor.

OpenStudy (anonymous):

M(0)=16

OpenStudy (anonymous):

So, using M(0)=16 and the updating function f(x)=0.75x+2 , I'd like to find the solution.

OpenStudy (anonymous):

Uh.. what is that \(f(5)\) for? That is neither the updating function nor is it an initial value...

OpenStudy (anonymous):

So, I should have typed \[M_5. \] = that whole mess.

OpenStudy (anonymous):

Well, let's solve the more general \(m_{t+1}=am_t+b\)... given \(m_0=c\), we have:$$m_1=ca+b\\m_2=a(ca+b)+b=ca^2+ab+b\\m_3=a(a^2c+ab+b)+b=a^3c+a^2b+ab+b\\...\\m_t=ca^t+b\sum_{k=0}^{n-1}a^k$$

OpenStudy (anonymous):

Now that we've figured that out, we can see our updating function is a special case where \(a=0.75\), \(b=2\), and \(c=16\). We can write our solution:$$(m(t)=16a^t+2\sum_{k=0}^{t-1}0.75^k$$

OpenStudy (anonymous):

Ok. That makes sense. I guess I couldn't figure out the way to express that solution. Because I saw the pattern going on, but couldn't figure out how to express it. Could you explain what the numbers around the summation symbol mean?

OpenStudy (anonymous):

Which numbers? The one out front is merely a coefficient... see how in \(m_3\) you have \(a^2b+ab+b=a^2b+a^1b+a^0b=b(a^2+a^1+a^0)\)?

OpenStudy (anonymous):

Yeah, so I see that pattern. I'm confused about how the sigma annotation, and how that would be computed. So, like, literally, how would I take \[M(t)=16(.75^t)+2\sum_{k=0}^{t-1}0.75^k\] and put it in a calculator?

OpenStudy (anonymous):

Do you have a graphing calculator? Most support summation.

OpenStudy (anonymous):

Yeah, I have a TI83.

OpenStudy (anonymous):

Sorry, I didn't really know you wanted a totally closed-form polynomial function... it's weird how similar difference and differential equations are, though! The solution should be: $$M(t)=8(0.75)^t+8$$

OpenStudy (anonymous):

http://mtl.math.uiuc.edu/node/87

OpenStudy (anonymous):

Hmmmm... Well, that equation looks a little better than the sigma one. Can you explain how you got it?

OpenStudy (anonymous):

http://www.cimt.plymouth.ac.uk/projects/mepres/alevel/discrete_ch14.pdf Sure! We're dealing with a first-order linear non-homogeneous difference equation with constant coefficients. The idea is that we can simplify a geometric progression rather easily. $$M(t)=16(0.75)^t+2\sum_{k=0}^{t-1}0.75^k\\\ \ \ \ \ \ \ =16(0.75)^t+2\left(\frac{1-0.75^t}{1-0.75}\right)\\\ \ \ \ \ \ \ =16(0.75)^t+8\left(1-0.75^t\right)\\\ \ \ \ \ \ \ =16(0.75)^t+8-8(0.75)^t\\\ \ \ \ \ \ \ =8(0.75)^t+8$$

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Geometric_progression#Geometric_series $$\sum_{k=0}^nr^k=\frac{1-r^{n+1}}{1-r}$$

OpenStudy (anonymous):

Sheesh, that is perfect. Thank you so much for all your help! I hadn't even heard of geometric progressions until just now. I don't know why my teacher didn't let us in on that one before assigning this. Thank you so much.

OpenStudy (anonymous):

No problem! Out of curiosity, what year of your university studies are you taking this in?

OpenStudy (anonymous):

Calculus 246 (It's Calculus 1 for Biology Majors.) My teacher might be assigning hard stuff, or I might just be behind. I took the prerequisites over a year ago... So that might be it. But thanks again.

OpenStudy (anonymous):

Neat! Are you a freshman? I'm a junior in high school so I'm trying to gauge how long I have to learn this stuff :-p

OpenStudy (anonymous):

Sheesh, nope. I'm a junior. I think a lot of the people in my class are freshmen straight out of high school though. I took Precal in my Junior Year in high school, Calculus in my Senior year, and then I took a few years off to travel before college. So when I jumped back in, I took Math 111 and Math 112 (College Algebra and Trig) and now I'm in Math 246. So, it's been a year since I took 112, and about 4-5 since I was deep into calculus. Are you planning on going straight through to college? If so, they'll probably let you take a placement test, and if you're math is fresh in your mind, you can probably save yourself a lot of tuition/time by getting into a more advanced math class. Good for you!

OpenStudy (anonymous):

(I am totally bummed that I didn't keep my math skills fresh and carry them straight through to college. If you can do that, you'll save yourself some frustration! :) )

OpenStudy (anonymous):

BTW, being taught by a high school junior simultaneously makes me feel dumb, and inspired by the internet/this website/high school kids. Keep up the good work! :)

OpenStudy (anonymous):

Yeah, I don't think I'll be taking any gaps... I'm taking as many math classes as I can before I graduate next year (Diff. EQ, lin. alg., etc.)! Don't consider yourself as being taught by me, I just tried to help you out a little... besides, I'm a math tutor at a local community college where I take courses!

OpenStudy (anonymous):

Smart plan. You're awesome. Good for you! :)